OpenStudy (anonymous):

evaluate the double integral by reversing the order of operation:

OpenStudy (anonymous):

\[\int\limits_{0}^{\sqrt{\pi}}\int\limits_{y}^{\sqrt{\pi}}Cos(x ^{2})dxdy\]

OpenStudy (anonymous):

i keep getting 0

OpenStudy (turingtest):

\[y\le x\le\sqrt\pi\]\[0\le y\le\sqrt\pi\]

OpenStudy (turingtest):

or is D on the other side...

OpenStudy (anonymous):

i don't know, all i was given is what i wrote

OpenStudy (turingtest):

No, I just need to figure it out...

OpenStudy (anonymous):

okay, can the answer to a double integral be a negative number?

OpenStudy (turingtest):

why not?

OpenStudy (anonymous):

idk, i thought it might be volume or something

OpenStudy (turingtest):

not, you are integrating a function over an area if the function is negative over that area, then the integral will be negative no magic there

OpenStudy (phi):

btw, 0 looks ok as an answer

OpenStudy (turingtest):

yeah, it is the region I drew with the D so\[y\le x\le\sqrt\pi\]\[0\le y\le\sqrt\pi\]implies\[x\le y\le\sqrt\pi\]\[0\le x\le\sqrt\pi\]so the integral is\[\large \int_{0}^{\sqrt\pi}\int_{x}^{\sqrt\pi}\cos(x^2)dydx\]um... did I mess up yet?